OpenStudy (anonymous):

Alex purchased a bedroom set for $2,276 using a six-month deferred payment plan with an interest rate of 23.49%. What is the balance after the deferment period if payments of $112 are made each month? $1,604.00 $1,884.74 $2,276.00 $2,556.74

jimthompson5910 (jim_thompson5910):

Hint: use A = P(1+r/n)^(n*t)

jimthompson5910 (jim_thompson5910):

another hint: the monthly payment figure they mention has no affect on the answer (it's put in there to throw you off)

OpenStudy (anonymous):

I know it's not $2556.74

OpenStudy (anonymous):

or C

jimthompson5910 (jim_thompson5910):

why not?

jimthompson5910 (jim_thompson5910):

A = P(1+r/n)^(n*t) A = 2276(1+0.2349/12)^(12*0.5) A = 2556.74

OpenStudy (anonymous):

Then I use the same formula as in the other question?

jimthompson5910 (jim_thompson5910):

no you're done

jimthompson5910 (jim_thompson5910):

that's the balance after 6 months

jimthompson5910 (jim_thompson5910):

so you were second-guessing yourself

OpenStudy (anonymous):

That's not the right answer

jimthompson5910 (jim_thompson5910):

it says 2556.74 is wrong?

OpenStudy (anonymous):

Yeah

jimthompson5910 (jim_thompson5910):

oooh let me try this

jimthompson5910 (jim_thompson5910):

oh wait nvm

jimthompson5910 (jim_thompson5910):

does it say which period after the deferment period ends?

jimthompson5910 (jim_thompson5910):

like "2 months after the deferment period"?

OpenStudy (anonymous):

Nope

jimthompson5910 (jim_thompson5910):

well the balance at the 6 month mark is 2556.74 if no payments are made during that 6 month window

jimthompson5910 (jim_thompson5910):

let's see what happens when payments of $112 are made instead

OpenStudy (anonymous):

What is the balance after the deferment period if payments of $112 are made each month?

OpenStudy (anonymous):

I think it's $1884.74 because if you multipy 112 by 6 months you get that answer

jimthompson5910 (jim_thompson5910):

one sec

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

hmm not sure why, but the closest I'm getting is 1850.99, but that's nowhere near $1884.74

jimthompson5910 (jim_thompson5910):

I wonder if it's implying that the interest is added after the payment is made

OpenStudy (anonymous):

Not so sure, I'll just go with $1,884.74

OpenStudy (anonymous):

Sakura purchased ski equipment for $1,248 using a six-month deferred payment plan. The interest rate after the introductory period is 23.79%. A down payment of $175 is required as well as a minimum monthly payment of $95. What is the balance at the beginning of the seventh month if only the minimum payment is made during the introductory period? A. $1,112.13 B. $637.13 Do you know which this one would be? I got $637.13 but i'm not to sure

jimthompson5910 (jim_thompson5910):

the down payment of 175 means only 1248-175 = 1,073 is financed

jimthompson5910 (jim_thompson5910):

one sec

jimthompson5910 (jim_thompson5910):

so that's the only two choices?

OpenStudy (anonymous):

No there were two more but those were wrong

jimthompson5910 (jim_thompson5910):

ok nvm that then

jimthompson5910 (jim_thompson5910):

I'm guessing the interest rate for the introductory period is not given or is it 0%?

OpenStudy (anonymous):

Not given

jimthompson5910 (jim_thompson5910):

hmm this is a tough one because you can choose not to pay during the deferment period (which I'm assuming is also the introductory period) but I'm trying various scenarios out and I'm not getting anything close I wish they would spell the process out clearer is there an example from the lesson we could use?

jimthompson5910 (jim_thompson5910):

if so please post it thanks

OpenStudy (anonymous):

No I looked through the lesson and everything was pretty crappy just like all these questions being given from this pretest

jimthompson5910 (jim_thompson5910):

hmm makes me wish I could read your book/lesson alongside you I'm ok at finance, but stuff like this is making me think otherwise lol

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

would you know how to do this one?

OpenStudy (anonymous):

Elliot is graduating from college in six months, but he will need a loan in the amount of $4,850 for his last semester. He may either receive an unsubsidized Stafford Loan with an interest rate of 6.8%, compounded monthly, or his parents may get a PLUS Loan with an interest rate of 7.8%, compounded monthly. The Stafford Loan has a grace period of six months from the time of graduation. Which loan will have a higher balance and by how much at the time of repayment? The PLUS Loan has a higher balance by $51.84. The Stafford Loan has a higher balance by $327.01. The Stafford Loan has a higher balance by $148.03. The PLUS Loan has a higher balance by $259.64.

OpenStudy (anonymous):

I'll show you what i did first

OpenStudy (anonymous):

Stafford Loan F = 4850*(1 + 0.068/12)^6 = 5017.25 PLUS Loan F = 4850*(1 + 0.078/12)^6 = 5042.25 Difference PL - SL = 5042.25 - 5017.25 = 28 but 28 isn't in any of the options

OpenStudy (anonymous):

F = L(1 + r/n)^nt F = Future Value L = Initial Loan r = Interest Rate in Decimal Form n = Number of Compounding Periods Per Year (intra-annual) t = Number of Years Loaned

jimthompson5910 (jim_thompson5910):

ok let me try it out

jimthompson5910 (jim_thompson5910):

Stafford Loan F = P(1+r/n)^(n*t) = 4850*(1 + 0.068/12)^(12*12/12) = 5,190.28 PLUS Loan F = P(1+r/n)^(n*t) = 4850*(1 + 0.078/12)^(12*6/12) = 5,042.25 Difference 5,190.28 - 5,042.25 = 148.03

jimthompson5910 (jim_thompson5910):

you forgot to do n*t in the exponent you just did t in the exponent

OpenStudy (anonymous):

oooooooh

jimthompson5910 (jim_thompson5910):

also, t is in years, so 6 months = 6/12 years 1 year = 12/12 months

OpenStudy (anonymous):

Alright, thanks so much!

jimthompson5910 (jim_thompson5910):

so The Stafford Loan has a higher balance by $148.03. yw

OpenStudy (anonymous):

Theres one more question I need help with, or are you annoyed of this?

jimthompson5910 (jim_thompson5910):

if i don't understand it, i get a bit frustrated lol but it's all in the learning process

jimthompson5910 (jim_thompson5910):

that last one wasn't so bad since it was definitely more straight-forward

jimthompson5910 (jim_thompson5910):

I wish I knew how to solve that second to last one though, which is why I wanted to see how the lesson would do it (to see an example)

OpenStudy (anonymous):

Shanelle purchased a dining room set for $2,620 using a 12-month deferred payment plan with an interest rate of 19.49%. She did not make any payments during the deferment period. What will the total cost of the dining room set be if she must pay off the dining room set within two years after the deferment period? $2,620.00 $3,864.00 $5,796.00 $3,178.82

OpenStudy (anonymous):

I think this is kinda like the question we could not figure out

jimthompson5910 (jim_thompson5910):

a bit but this one is more clear though, it says "She did not make any payments during the deferment period"

jimthompson5910 (jim_thompson5910):

A = P(1+r/n)^(n*t) A = 2620(1+0.1949/12)^(12*1) A = 3178.82 note: this is NOT the answer and a lot of people think it is (which is why this trap is thrown in here). This value is used to find the final answer.

jimthompson5910 (jim_thompson5910):

3178.82 is the balance after the 12 month deferment period let's use this to find the monthly payment Use the formula P = L((r/n)*(1 + r/n)^(n*t))/((1 + r/n)^(n*t) - 1) where, P = monthly payment L = total amount loaned or amortized r = annual interest rate (APR) n = number of times interest is compounded per year (compounding frequency) t = time in years In this case, P = unknown (we're solving for this) L = 3178.82 r = 0.1949 (19.49% = 19.49/100 = 0.1949) n = 12 t = 2 P = L((r/n)*(1 + r/n)^(n*t))/((1 + r/n)^(n*t) - 1) P = 3178.82((0.1949/12)*(1 + 0.1949/12)^(12*2))/((1 + 0.1949/12)^(12*2) - 1) P = 160.99771204627 P = 161 So we know Payment per period: P = $161 Number of periods: n*t = 12*2 = 24 Total Amount Paid = (Payment per period)*(Number of periods) = ($161)*(24) = $3864 Total Cost is $3864.00

OpenStudy (anonymous):

Thank you so much! I submitted it now and the Sakura one was 637.13

jimthompson5910 (jim_thompson5910):

thanks, I'll have to go back over that and think how they got that

jimthompson5910 (jim_thompson5910):

oh what was the answer to "Alex purchased a bedroom set for..." I don't think we got that either...hmm

OpenStudy (anonymous):

It was $1,884.74

jimthompson5910 (jim_thompson5910):

Ok I figured it out. Here's how it works. The rules are that if you pay off the entire balance (this is a big IF), then NO interest will apply. So if you manage to pay off the entire balance of $2,276 within 6 months (the deferred interest period), then you will be charged NO interest at all. However, companies know very well that the majority of the people will not be able to pay off the entire balance within 6 months. So this is when they retroactively charge interest to make a lot of money. So Alex could have made payments to fully pay off the $2,276 debt within 6 months...BUT...Alex made payments of $112 for 6 months, which means he really only paid 6*112 = 672 dollars toward the balance and he's nowhere close to paying off the entire balance of $2,276 So here's what happens a) Alex did make 6 payments of $112 or $672 total over 6 months. So subtract this from the initial balance to get: 2,276 - 672 = 1604. Notice how the balance is not $0. So the balance was not paid for in full. If it was $0, then Alex can walk away without having to make any more payments. It's not $0, so we move onto b) b) Because the balance was NOT paid in full by the end of the 6 month deferment period, this means that interest is applied for every month in the 6 month deferment period. All of this is applied to the initial balance and payments are not factored in (yet). So, P*(1+r/n)^(n*t) = 2276*(1+0.2349/12)^(12*6/12) = 2,556.74 c) The balance is now 2,556.74 dollars. This would be the final answer if Alex did not make any payments at all during this 6 month period. However, alex did make monthly payments of $112 and he paid the company $672 so far. So that is subtracted from the balance to get 2,556.74 - 672 = 1,884.74 So that explains why the remaining balance after the 6 month deferment period is $1,884.74 The same basic steps are applied to the problem with Sakura as well.

OpenStudy (anonymous):

Yeah, that's what I did

OpenStudy (anonymous):

For the Alex one

jimthompson5910 (jim_thompson5910):

ok great, glad you figured it out

OpenStudy (anonymous):

Hey sorry to bother you but theres a question just like the Alex one Farrah installed a new pool for $14,730 using a 12-month deferred payment plan with an interest rate of 19.33%. What is the balance after the deferment period if payments of $527 are made each month? $14,370.00 $17,843.62 $8,046.00 $11,519.62 My brother tried to do this with me and he got $8046.00 whereas I got $11,519.62 I did exactly what you said for the Alex question like 14730(1+0.1933/12)^(12*12/12) = 17843.62 THEN I took 527*12months and got 6324 Subtracted 17843.62 - 6324 = 11519.62 I'm pretty confident with my answer but my brother is telling me I'm wrong?

jimthompson5910 (jim_thompson5910):

your brother would be right if the interest wasn't applied retroactively, but it is

jimthompson5910 (jim_thompson5910):

$11,519.62 is the correct answer

jimthompson5910 (jim_thompson5910):

it would be nice if it was as simple as saying initial balance - (# of months)*(payment per month) 14730-12*527 8,406 but it's not that simple and that trap is thrown in there to catch students off guard

OpenStudy (anonymous):

Ah okay, thanks :) again!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

Garrett is graduating from college in twelve months, but he will need a loan in the amount of $6,785 for his last two semesters. He may either receive an unsubsidized Stafford Loan with an interest rate of 6.8%, compounded monthly, or his parents may get a PLUS Loan with an interest rate of 7.8%, compounded monthly. The Stafford Loan has a grace period of six months from the time of graduation. Which loan will have a higher balance at the time of repayment and by how much? The PLUS Loan has a higher balance by $72.54. The PLUS Loan has a higher balance by $112.83. The Stafford Loan has a higher balance by $177.86. The Stafford Loan has a higher balance by $250.40 Is C the right answer?

jimthompson5910 (jim_thompson5910):

Stafford Loan A = P(1+r/n)^(n*t) A = 6785(1+0.068/12)^(12*1.5) A = 7511.43390604005 A = 7511.43 ------------------------------- PLUS Loan A = P(1+r/n)^(n*t) A = 6785(1+0.078/12)^(12*1) A = 7333.56596333002 A = 7333.57 ------------------------------- The Stafford Loan has the higher balance Difference: 7511.43 - 7333.57 = 177.86 So C is definitely the correct answer